1/1

Marie von Ebner-Eschenbach,

A little song

tradukita de Frank Petersohn

A little song, How does it go? How dear to me It speaks, but lo, Why is it so? Relate, do!

Hark, it contains A pleasant ring, Comforting strains, Some words to sing And someones entire soul too.

Traduko de la Germana poemo “Ein kleines Lied” de Marie von Ebner-Eschenbach (∗1830 – †1916) en la Anglan de Frank Petersohn.

Arg-46-1464 (2012-12-27 12:41:06)

Tiu ĉi poem-angligo troviĝas en http: // ingeb. org/ Lieder/ einklein. html .

http://www.poezio.net/version?poem-id=46&version-id=,1464